OpenStudy (anonymous):

If the hypotenuse of a right triangle is 2√3 units long and one of the legs is √3 units long, then how long is the other leg? A.) 3 B.) 2√3 C.) √3

OpenStudy (anonymous):

isnt it the same tho? two legs are the same

OpenStudy (anonymous):

so my guess is c?

OpenStudy (anonymous):

This will be a 30-60-90 triangle. And artie is wrong

OpenStudy (anonymous):

whats the answer?

OpenStudy (anonymous):

The relation between the hypotenuse and small leg is that the small leg is 1/2 of the hypotenuse. So the hypotenuse is \(2\sqrt{3}\) and the small leg is \(\sqrt{3}\). The leg we are looking for is the long leg, which is \[\large small~leg\sqrt{3}\]